### v1

@@ 5.2.2 Method limitations
Currently, the depth of our architecture is limited due to vanishing gradient as we still lack some widely adopted architectural elements that mitigate this effect, as exposed in Appendix A.8.1. Even considering our limited accessible architecture space, and despite exploring tens of variations, we might have missed some specific architecture construction that would outperform our YOLO-CIANNA-ref model. It goes the same way regarding our selection of hyper-parameter values, especially considering their number and their strong interdependence.
@@
An obvious limitation of our YOLO-CIANNA method is that it is supervised. In this regard, we already discussed how the training area and catalog representativity affect the capability of our method. We also presented the importance of the purity and completeness of the training sample when designing our selection function (Sect. 3.4), which we consider still unsatisfying despite several attempts to improve it. Consequently, the trained model is constrained to a very specific instrumental setup, and generalization toward other datasets is likely to be difficult (Sect. 5.3). There are ways to force the network to be more independent of the exact training dataset. For example, it can be forced to base its decision on relative contrasts and morphologies rather than the absolute signal, but the absolute source flux would have to be retrieved differently. We could also have a single network trained on images from various instruments and configurations and provide the instrumental setup as an external input. This way, the common features between similar images would be mutualized into a single model.
@@
Then, we have limitations that are linked to the construction of our method. The most striking is that we predict bounding boxes to detect mostly elliptical objects that typically fade with the distance to the center. In practice, the network has to fit luminosity profiles with a significant part that can be below the noise level to identify the objects properly. This task is inherently different than searching for the box that contains all the visible pixels of a given object, which justifies several adaptations we made in both the method itself and the network backbone architecture. This specific property of astronomical source detection explains why most source detection methods are based on segmentation. The main issue with segmentation methods is that they are compute-intensive. In this regard, bounding-box-based detection methods are more efficient in discriminating what part of the input signal is relevant to the task to perform. From this, we consider that the best approach would be to apply segmentation methods on restrained regions around the sources detected with another method like ours.
@@
Another structural limitation is that we must know in advance the density of sources that we expect to detect and at which typical size and resolution. Starting with YOLO-V3, the classical implementation adds the possibility of having detection at different reduction factors using the same network, which slightly elevates this issue. In addition, grid-based detection can bias the estimated confidence of objects near grid element edges and construct an artificial discretization of the detection space. A possible solution would be to have overlapping grid cells, which would complexify the association process and the post-process filtering. As a long-term objective, we would like to modify our method to recursively extract boxes of different sizes until it considers that nothing is left in the detection area, which is likely to be feasible using transformer principles (e.g., Carion et al., 2020; Zhang et al., 2021; Fang et al., 2021). By doing this, we could remove the need to define the number of required detection units.
@@
In our method, we already strongly modify the association process in comparison to the classical YOLO approach (Sect. 2.6 and Appendix A.3). By using a prediction-aware association process, we added more information about the current network confidence to guide the training. However, we limited our matching score metric to geometrical properties. Another approach would be to include some of the additional predicted properties. For example, we could include the predicted flux and position angle of a source to select the best current detection in the association process, which would be more similar to classical astronomical cross-matching criteria.
@@ 5.3 Generalization to real observational data
In this paper, we evaluated and discussed the performances of YOLO-CIANNA on a simulated dataset. We will now have to evaluate its capability over real observational data. The only strong requirement for the method is to have a training sample representative of the dataset to which it will be applied. Several approaches could generate a confident target dataset compatible with an application over observational data.
@@
One approach would be simulations since they provide complete and unambiguous target catalogs. The difficulty is that it must be as realistic as possible regarding source shapes and properties. But more importantly, it must be realistic regarding all the elements the detector must ignore. This includes possible diffuse signals, characteristic instrumental noise, or artifacts of various kinds and origins. Overall, the instrumental model is often more critical regarding supervised detector performances than the target catalog, which is an identified weakness of the SDC1 dataset (Sect. 3.2). Any lack of realism in the training sample would result in missed interesting objects, false detections, or unpredictable behaviors due to unconstrained contexts. While the SDC1 dataset can be a great way to train supervised approaches before applying them to similar radio astronomical data, there is much more value in the underlying T-RECS simulation that was used to construct it (Bonaldi et al., 2019). Such simulation could generate simulated observations for various radio telescopes on which dedicated detectors could be trained. We also note that much effort has been put into the realism of the instrumental data pipeline for the simulated data used in the latest SDC3. These improvements could be used to construct a new simulated continuum dataset. In practice, even if a detector trained solely on simulated data could be directly applied to observational data, it would be more realistic to consider the simulated images as pretraining datasets and always have a fine-tuning step over some real images.
@@
Another approach would be to use solely observed data to define a training sample for a given instrument. While realism is assured, it is more difficult to define the target examples. One classical approach in computer vision is to rely on human labeling through crowd science. While it was already extensively used for astronomical datasets and noticeably for galaxy identification (e.g., Lintott et al., 2008; Banfield et al., 2015), it presents several difficulties. First, analyzing this type of dataset usually requires more expert knowledge than tagging everyday-life images. Then, due to the source density, the diversity of morphologies, and the variety of contexts, a large crowd-science campaign would be required to obtain exploitable results. This limitation explains why some studies only call for visual inspection of litigious cases after automated detection. The more robust and classical alternative would be to obtain strong observational confirmations for a limited number of examples using other instruments or simply with more integration time on the same instrument. The difficulty is that instrument time is expensive and difficult to obtain. A possible solution would be to observe only a limited amount of source candidates obtained from a detector trained in another way to refine its training sample.
@@
In practice, we have already attempted to apply our YOLO-CIANNA-ref model to surveys from SKA precursor instruments. We noticeably explored the respective closest frequency from the LOFAR (LOw-Frequency ARray) Two-metre Sky Survey (LoTSS, Shimwell et al., 2022), and the Rapid ASKAP (Australian Square Kilometre Array Pathfinder) Continuum Survey (RACS, McConnell et al., 2020). Performing a direct application of our SDC1-trained model over images from instruments with a different resolution and sensitivity requires adapting the image dynamic and the object’s apparent size to be similar to what was seen during training. Then, the detector sensitivity can be adjusted through the objectness thresholds to achieve the desired completeness or purity. Our preliminary results with this approach are encouraging for both surveys, the main drawback being a too high amount of false detection around bright sources that produce radial artifacts. Still, we are confident that our method will be capable of understanding such context through complementary training using observational examples.
@@
As exposed in Sect. 3.9, our method implementation allows for a very compute-efficient source-detector. Our method is perfectly suited for real-time object detection, which could be used either to perform low-level analysis and flagging inside data-heavy pipelines or as an integrated module for added-value data visualization and exploration. We are currently exploring how to deploy our method as a standardized service using virtual observatory tools to interact with image viewers through an API. It would allow users to perform detection and characterization on the fly from a list of pre-trained models with the possibility of adjusting several prediction-related hyper-parameters. We are developing experimental support for these functionalities in the YaFITS viewer (Salome et al., 2021).
@@ 6 Conclusion
In this paper, we presented a new deep-learning source detection and characterization method called YOLO-CIANNA that was inspired by the widely adopted YOLO regression-based object detector. We used the SKAO SDC1 dataset as a benchmark for evaluating our method capability on simulated continuum radio images that should resemble upcoming SKA data products.
@@
We highlighted that astronomical data presents specific properties that challenge the classical design of deep-learning detection methods. Typical astronomical images representing large fields with high sensitivity have a huge dynamical range and are crowded with small objects of a few pixels that can be blended. We introduced the new elements of our custom method that account for these specificities, like using a prediction-aware association function during training and putting the focus on the central position prediction. We added to our method the capability of predicting an arbitrary number of additional parameters, allowing it to characterize the detected astronomical sources while preserving a single prediction stage. We also expanded our loss to include more conditional contributions, resulting in a cascading loss process that progressively guides the detector expressivity toward a more complex and complete problem as it gets better at solving the task. In the dedicated Appendix A.3, we provided more details about our custom association process and highlighted the differences with the classical YOLO implementation. Our method achieves the same level of performance as the classical YOLO implementation for classical computer vision datasets using a given network backbone.
@@
We identified that some specificities of the challenge design required special attention. Deep-learning methods being extremely sensitive to the completeness and purity of the training sample, we presented a selection function that filters extract the detectable sources from the complete reference catalog. In its default configuration, the SDC1 training area does not represent the full image due to the shape of the primary beam response. We presented an approach that uses the accessible full image to enrich the training sample with noise-only fields. We demonstrated that this approach allows our detector to interpolate between these two training regimes and achieve good results at the scale of the full image. In a dedicated Appendix C, we propose an alternative training area that represents a radius of the full image, which proved to solve the representativity issue.
@@
We introduced a custom network architecture backbone for our method that we optimized regarding the specificities of astronomical images while preserving the relevant structural element of a typical YOLO darknet-19 architecture. In a dedicated Appendix B, we evaluate the performance of the default darknet-19 architecture backbone on the SDC1 task using our adapted detection formalism. This test showed that architectures designed for everyday-life images are not suited for astronomical data and that our custom architecture significantly outperforms it. We also disprove the classical assumptions that low-level features trained on everyday-life images would be good candidates for domain adaptation toward astronomical data.
@@
For performing the prediction over the large $32768{\times}32768$ input image of the SDC1, we established a prediction pipeline that decompose it into overlapping patches used as individual input for our detector. By combining our method with a fully convolutional network backbone, it is possible to use any input image size for the patches and obtain roughly the same results, except for edges and receptive field effects. We presented a complete post-processing approach that removes multiple detections while preserving blended sources that are confidently detected and that account for the overlapping areas between patches.
@@
In its current setup, our YOLO-CIANNA detector strongly outperformed the other methods that were applied to the SDC1 dataset. We achieved a +139% score improvement in the challenge metric compared to the team that reached first place during the challenge. Even in comparison with another post-challenge result, we still achieve a score that is higher by 61%. Our reference catalog has a detection purity of more than 94% while detecting 40 to 60% more sources than the other top-score methods. We also demonstrated that it is possible to constrain the purity regime of a given trained model in post-processing. This capability was used to produce a catalog with 99% purity that still detects 10 to 30% more sources than the other top-score methods. Our YOLO-CIANAN detector presents excellent completeness and purity over a large apparent flux range. Our detector is especially efficient at low SNR regimes, and we are likely to have reached a structural limit on this aspect. Regarding source characterization, our method also significantly outperforms all the other submissions.
@@
We are working on applying our YOLO-CIANNA detector to observational data from the LOFAR and RACS instruments and are confident that it will be a great tool for analyzing upcoming SKA science data products. Our method is also very compute efficient, allowing us to reach real-time detection capability that could be deployed as a service and integrated into astronomical image viewers, especially in the context of the future SKA SRC.
@@
For reproducibility purposes, the training and prediction scripts that allow us to construct our YOLO-CIANNA-ref result are provided alongside this paper in the form of an example directory dedicated to the SDC1 in our CIANNA repository. We also provide simplified notebooks for visualizing the model results and making discovering the method itself easier. The exact scripts are archived with the CIANNA V1.0 release, while the YOLO-CIANNA-ref model is archived with an independent DOI and linked as a data product of CIANNA.
@@ Acknowledgements.
We acknowledge the support of the MINERVA project for the Paris Observatory, including its material and financial support. We acknowledge the Paris Observatory DIO shared computing resources and their handling of the MINERVA dedicated compute resources. We acknowledge the computing resources from the GENCI (Jean-Zay GPU) accessed in the context of MINERVA’s team participation in the SDC2. DC acknowledges its PLS fellowship. We highlight the great work done by SKAO in organizing all the SDCs and thank them for allowing post-challenge access to the corresponding datasets.
@@ Appendix A Method details and comparison to the classical YOLO
The classical YOLO method has been specifically designed for detecting objects in everyday life images. These usually respect some implicit properties like objects with sharp features, objects obscuring each other, images being not too noisy and with a limited dynamic range, images being relatively not crowded, and detectable objects being relatively large in pixel size. In contrast, astronomical images can have a considerable dynamic range, be noisy or have strong and complex instrumental artifacts, be crowded with objects that might be as small as a few pixels, have blending between objects, and have object types or classes that are very degenerate. Some of these difficulties were addressed at various places in Sects. 2 and  3. In this section, we expose some specific differences between our method and the classical YOLO implementation that should make it more suitable for astronomical source detection. We focus on improving the capability to detect multiple similar objects in a given grid cell and the cascading loss principle we implemented to balance the network expressivity. Finally, to verify that our modified method still performs well for classical object detection, we apply it to some standard computer vision datasets and compare our results to the classical YOLO-V2.

### v2

@@ 5.2.2 Method limitations
Currently, the depth of our network backbone architecture is limited due to vanishing-gradient. We identified that group normalization and other variants known to mitigate this issue are detrimental to our detection and characterization performances with our current setup. The best solution would be to use residual or skip connections (He et al., 2016), but those are not yet available in our generic CIANNA framework. The absence of these elements also prevents us from adding an efficient multi-scale detection in our YOLO-CIANNA method, as is the case in most modern object detectors. Even considering our limited accessible architectural space, and despite exploring hundreds of variations, we might have missed some specific architectures that would outperform our current YOLO-CIANNA model. It goes the same way regarding our selection of hyperparameter values, especially considering their number and strong interdependence.
@@
Another obvious limitation of our YOLO-CIANNA method is that it is fully supervised. In this regard, we already discussed how the training area and catalog representativity affect the capability of our method. We also presented the importance of the purity and completeness of the training sample when designing our selection function (Sect. 3.4), which could still be improved. Consequently, the trained model is constrained to a specific instrumental setup, and generalization toward other datasets is likely to be difficult (Sect. 5.3). There are ways to force the detector to be more independent of the exact training dataset. For example, it can be incentivized to base its decision on relative contrasts and morphologies rather than the absolute signal. Still, the absolute source flux would have to be retrieved differently. We could also have a single network trained on images from various instruments and configurations and provide the instrumental setup as an external input. This way, the shared features between similar images would be mutualized into a single model.
@@
Then, we have limitations linked to the construction of the method itself. The most striking is that we predict bounding boxes to detect mostly elliptical objects that typically fade with the distance to the center. In practice, the detector has to fit luminosity profiles, with a significant part that can be below the noise level, to identify the objects correctly. This task is inherently different than searching for the box that contains all the visible pixels of a given object, which justifies several design choices in the method itself and the network backbone architecture. This specific property of astronomical source detection explains why most source detection methods are based on segmentation. The main issue with segmentation methods is that they are computationally intensive. In this regard, bounding-box-based detection methods are more efficient in discriminating what part of the input signal is relevant to the specific task. From this, we consider that the best approach would be to use segmentation methods on small regions around sources detected with a method like ours.
@@
Another structural limitation is that we must know in advance the density of sources we expect to detect and at which typical size and resolution. Several modern deep-learning object detectors implement multi-scale detection capabilities. This is usually done by placing detection heads periodically in the network between spatial dimension reduction operations or adding skip connect at similar places toward the single final detection head. This way, the network can access information from small scales before they are compressed into larger-scale patterns (e.g., Lin et al., 2017; Redmon & Farhadi, 2018). In addition, grid-based detection can bias the estimated confidence of objects near grid element edges and construct an artificial discretization of the detection space. A possible solution would be to have overlapping grid cells, which would complexify the association process and the post-process filtering. As a long-term objective, we would like to modify our method to recursively extract boxes of different sizes until nothing is left in the detection area, which should be feasible using transformers (e.g., Carion et al., 2020; Zhang et al., 2021; Fang et al., 2021). It means we could remove the need to define the number of detection units.
@@
By using a prediction-aware association process, our method exploits information about the current detector confidence to guide the training. However, we limited our matching score metric to geometrical properties. Another approach would be to allow some of the additional predicted properties to constrain the association. For example, we could include the predicted flux and position angle of a source to select the best current detection in the association process, which would be more similar to classical astronomical cross-matching criteria. As we observed that predicting correlated extra-parameters improves the detection-only score, we are confident that this addition would accelerate and stabilize the training process of our detector and likely result in higher detection scores. We note that this would only affect the association phase and that the resulting trained model could still be applied to a detection-only task with no prediction of the source flux or other extra parameters.
@@ 5.3 Generalization to real observational data
In this paper, we evaluated and discussed the performances of YOLO-CIANNA on a simulated dataset. We still have to evaluate its capability over real observational data. The only strict requirement for the method is to have a training sample representative of the dataset to which it will be applied. Several approaches could be employed to generate a confident target dataset compatible with an application over observational data.
@@
One approach would be simulations since they provide complete and unambiguous target catalogs. The difficulty is that it must be as realistic as possible regarding source shapes and properties. But more importantly, it must be realistic regarding instrumental properties and background signals. This includes possible diffuse structures, characteristic instrumental noise, or artifacts of various kinds and origins. Overall, the instrumental model is often more critical regarding supervised detector performances than the target catalog, which is an identified weakness of the SDC1 dataset (Sect. 3.2). Any lack of realism in the training sample would result in missed interesting objects, false detections, or unpredictable behaviors due to unconstrained contexts. While the SDC1 dataset could be used to train supervised methods before applying them to similar radio astronomical data, there is much more value in the underlying T-RECS simulation that was used to construct it (Bonaldi et al., 2019). Such simulation could generate mock observations for various radio telescopes on which dedicated detectors can be trained. We also note that much effort has been put into the realism of the instrumental data pipeline for the simulated data used in the latest SDC3. These improvements could be used to construct a new simulated continuum dataset. In practice, even if a detector trained solely on simulated data could be directly applied to observational data, it is often better to consider the simulated images as pretraining data and have a fine-tuning step over some real data, which is a case of transfer learning.
@@
Another approach would be to use solely observed data to define a training sample for a given instrument. While realism is assured, it is more difficult to define the target list. One classical approach in computer vision is to rely on human labeling through crowd science. While it was already extensively used for astronomical datasets and notably for galaxy identification (e.g., Lintott et al., 2008; Banfield et al., 2015), it presents several difficulties. First, analyzing astronomical datasets requires more expert knowledge than tagging everyday-life images. Then, due to the survey size, source density, diversity of morphologies, and variety of contexts, a large crowd-science campaign would be required to obtain exploitable results. This limitation explains why some studies only call for visual inspection of litigious cases after automated detection. The more robust and classical alternative would be to use strong observational confirmations for a limited number of examples. A typical example would be to observe a small region of a large survey using the same instrument but with a much longer integration time. This way, a classical detection method can be employed to construct a robust training catalog for a deep learning detector that takes input images at the typical survey lower integration time. Such a detector is expected to produce better results than a classical method at the application integration time and can then be used over the whole survey. However, instrument time is expensive and difficult to obtain. A possible solution would be to get strong confirmations only for a limited number of source candidates obtained with a detector trained from a less confident source catalog (see Appendix C.2).
@@
We tried to apply our SDC1-trained model to surveys from SKA precursor instruments. We notably explored the respective closest frequency from the LOFAR (LOw-Frequency ARray) Two-metre Sky Survey (LoTSS, Shimwell et al., 2022), and the Rapid ASKAP (Australian Square Kilometre Array Pathfinder) Continuum Survey (RACS, McConnell et al., 2020). Performing a direct prediction over images from instruments with a different resolution and sensitivity requires adapting the image dynamic range and the apparent size of the objects to be similar to what was seen during training. Then, the detector sensitivity can be adjusted through the objectness thresholds to achieve the desired completeness or purity. Our preliminary results with this approach are encouraging for both surveys, the main drawback being that there is a high amount of false detection around bright sources that produce radial artifacts. Still, we are confident that our method will be capable of understanding such context through complementary training using observational examples.
@@
As exposed in Sect. 3.9, our method can be used to produce a computationally efficient source-detector, making it well suited for real-time object detection. It could be used to perform low-level analysis and flagging inside data-heavy pipelines or as an integrated module for added-value data visualization. We are currently exploring how to deploy our method as a standardized service using virtual observatory tools to interact with image viewers through an API. It would allow users to perform detection and characterization on the fly from a list of pretrained models with the possibility of adjusting several prediction-related hyperparameters. We are developing experimental support for these functionalities in the YaFITS viewer (Salome et al., 2021).
@@ 6 Conclusion
In this paper, we present a new deep-learning source detection and characterization method called YOLO-CIANNA, which takes inspiration from the widely adopted YOLO regression-based object detector. We use the SKAO SDC1 dataset as a benchmark for evaluating the capability of our method when applied to simulated continuum radio images that should resemble upcoming SKA data products.
@@
We highlight that astronomical data have specific properties that challenge the classical design of deep-learning detection methods. Typical astronomical images representing wide fields with high sensitivity have a huge dynamic range and are crowded with small objects of a few pixels, which can be blended. We present several low-level functionalities and discuss method design choices introduced to handle these specificities, like using a conditional prediction-aware association function during training and putting the focus on the central position prediction. This results in a cascading loss process that progressively guides the detector expressivity toward a more complex and complete problem as it gets better at solving the task. Our method is also capable of predicting an arbitrary number of additional parameters, allowing it to characterize the detected astronomical sources while preserving a single prediction stage. In the dedicated Appendix A.3, we provide more details about our custom association process and highlight the main differences from the classical YOLO method. We also show that our method achieves the same level of performance as YOLO-V2 for classical computer vision datasets using a given network backbone.
@@
Some specificities of the SDC1 challenge design require special attention. We present a selection function that extracts the detectable sources from the complete catalog. In its default configuration, the SDC1 training area does not represent the full image because of the shape of the primary beam response. We present an approach that uses the accessible full image to enrich the training sample with noise-only fields. We demonstrate that this approach allows our detector to interpolate between these two regimes and to achieve good results at the scale of the full image. In the dedicated Appendix C, we propose an alternative training area that represents a beam radius over the image field, and allows us to solve the representativity issue.
@@
We introduce a custom network architecture backbone for our method that we have optimized regarding the specificities of astronomical images. In the dedicated Appendix B, we evaluate the performance of a classical darknet-19 architecture backbone on the SDC1 task using our adapted detection formalism. This test shows that architectures designed for everyday-life images are not suited to astronomical images and that our custom architecture significantly outperforms it. We also disprove the classical assumptions that low-level features trained on everyday-life images are good candidates for domain adaptation toward astronomical images.
@@
To perform the prediction over the large $32768{\times}32768$ input image of the SDC1, we establish a prediction pipeline that decomposes it into overlapping patches used as individual input for our detector. We present a complete post-processing approach that removes multiple detections while preserving blended sources that are confidently detected and that account for the overlapping areas between patches.
@@
In its current setup, our YOLO-CIANNA detector strongly outperforms other methods applied to the SDC1 dataset. We achieved a +139% score improvement in the SDC1 scorer metric compared to the team that reached first place during the actual SDC1 challenge. Even in comparison with another post-challenge result, we still achieve a score that is higher by 61%. Our catalog has a detection purity of more than 94% while detecting 40 to 60% more sources than the other top-score methods. We also demonstrated that it is possible to constrain the purity regime of a given trained model in post-processing. This capability was used to produce a catalog with 99% purity that still detects 10 to 30% more sources than the other top-score methods. Our detector is especially efficient in low-S/N regimes, and we are likely to have reached a structural limit on this aspect. Regarding source characterization, our method also significantly outperforms all the other submissions. The corresponding source catalogs are archived at 10.5281/zenodo.13141772.
@@
We are working on applying our YOLO-CIANNA detector to observational data from the LOFAR and RACS instruments and expect it to be a great tool for analyzing upcoming SKA science data products. Our method is also computationally efficient, allowing us to reach real-time detection capability. As a result, it could be deployed as a service and integrated into astronomical image viewers, especially in the context of the future SKA science regional centers.
@@
For reproducibility purposes, the training and prediction scripts that allowed us to construct our YOLO-CIANNA result are provided alongside this paper in the form of an example directory dedicated to the SDC1 in our CIANNA repository. We also provide simplified notebooks for visualizing the model results and to make it easier to explore the method itself. The exact scripts are archived with the CIANNA V-1.0 release (10.5281/zenodo.12806325), while our reference YOLO-CIANNA model for the SDC1 is archived at 10.5281/zenodo.12801421.
@@ Acknowledgements.
We acknowledge the support of the MINERVA project from the Paris Observatory, including its material and financial support. We acknowledge the Paris Observatory DIO shared computing resources and their handling of the MINERVA dedicated compute resources. We acknowledge the computing resources from the GENCI (Jean-Zay GPU partition) provided to team MINERVA in the context of the SDC2. DC acknowledges the MINERVA project for funding two years of post-doctoral contract. DC acknowledges its PSL fellowship as part of the MCA-IA program. We highlight the great work done by SKAO in organizing all the SDCs and thank them for allowing post-challenge access to the corresponding datasets.
@@ Appendix A Method details and comparison to the classical YOLO
Classical object detection methods have usually been designed explicitly to detect objects in everyday life images. These respect some implicit properties like objects with sharp features, objects obscuring each other, images being not too noisy and with a limited dynamic range, images being relatively not crowded, and detectable objects being relatively large in pixel size. In contrast, as stated in Sect. 1, astronomical images can have a dynamic range of several orders of magnitude, be noisy, have complex instrumental artifacts, be crowded with objects that might be as small as a few pixels, have blending between objects, and have object types or classes that are very degenerate. Some of these difficulties were highlighted at multiple places in Sects. 2 and  3. In this appendix section, we expose some specificities of our YOLO-CIANNA method that contrast with the classical YOLO method, making it more suitable for astronomical source detection. We first discuss how to achieve simultaneous detection of multiple objects per grid cell and the cascading loss principle we implemented to balance the network expressivity. Finally, to verify that our method performs well for classical object detection, we apply it to some standard computer vision datasets and compare our results to the classical YOLO-V2 using an almost identical CNN backbone.